A light bulb manufacturer claims that the mean life of a certain type of light bulb is 750 hours. If a random sample of 36 light bulbs has a mean life of 725 hours with a standard deviation of 60 hours. Use a=0.05a. State the null and alternative hypotheses.
b. State the Type I and Type II errors.
c. Find the critical value. Do you have enough evidence to reject the manufacturer’s claim?
d. Find the p-value.
e. Construct a 95% confidence interval for the population mean.
Answer #1

Identify the null hypothesis and alternative hypothesis:    

H_0: \mu\geq 720 \ (claim) (Bulbs \ last\ at\ least \ 720\ hours)

H  0   :μ≥720 (claim)(Bulbs last at least 720 hours) H_1: \mu<720 \ (Bulbs\ last\ less\ than\ 720\ hours)\ \ \ \ \ \ \ \ \ \

 Identify the critical value(s):  Left-tailed test    z_\alpha=z_{0.05}=-1.645\approx-1.65z  α   =z  0.05   =−1.645≈−1.65    The critical region will be to the left of –1.65.    Identify the standardized test statistic    

z={\bar{x}-\mu \over \sigma/\sqrt{n} }={712.8-720 \over 62/\sqrt{51} }\approx-0.83z=  σ/  n    x ˉ  −μ   =  62/  51    712.8−720   ≈−0.83 

Decide whether to reject or fail to reject the null hypothesis and interpret the decision in the context of the original claim  A one-tailed test (right or left) indicates that the null hypothesis should be rejected when the test value is in the critical region on one side of the mean.  The rejection zone is below -1.65 at the 0.05 level.  Fail to reject H_0.H  0. There is not sufficient evidence to reject the claim that means bulb life is at least 720 hours.    

A type I error occurs if one rejects the null hypothesis when it is true.  

Type I Error: Incorrectly claim light bulbs last less than 720 hours    A type II error occurs if one does not reject the null hypothesis when it is false.  

Type II Error: Fail to detect that light bulbs last less than 720 hours

answered by: anonymous